# On Robust hypothesis testing with respect to Hellinger distance

### v1

@@ 1-B Robustness to imperfect modelling
$\mathcal{H}_{0}$ $:\text{Pois}(\lambda_{1})$
$\mathcal{H}_{1}$ $:\text{Pois}(\lambda_{2})$
@@
$H_{0}$ $:\gamma d(p,p_{1})\leq d(p,p_{1})$
$H_{1}$ $:d(p,p_{1})\geq\gamma d(p,p_{1}).$
If $p$ is not in $H_{0}$ or in $H_{1}$, then the test can make any decision. The error is defined to be the maximum of type-I and type-II errors. We call a test $\gamma$-robust if it uses finitely many samples and keeps the probability of error below some fixed constant $\delta<\frac{1}{2}$ for any choice of $p_{1}$, $p_{2}$, and pp. Let $\gamma^{*}$ be the smallest $\gamma$ for which a $\gamma$-robust test exists. We call $\gamma^{*}$ the optimal slack factor for the robust testing problem. Previous works have tried to characterize $\gamma^{*}$ for different distance measures $d(.,.)$. The $\ell_{p}$-norm of a function $f:\mathcal{X}\rightarrow\mathbb{R}$ is given by
@@
where $A=\{x^{n}:p_{1}(x^{n})>p_{2}(x^{n})\}$ and $\mu(A)=\frac{1}{n}\sum_{i=1}^{n}1[X_{i}\in A]$, that is, the fraction of observed samples which fall in the set AA. The lower bound ($\gamma^{*}\geq 3$) was given in the work of Bousquets, Kane, and Moran [7]. The Hellinger distance between $p_{1}$ and $p_{2}$ is given by
@@ 3 Upper Bound on $\gamma^{*}$
What kinds of tests are $\gamma$-robust for some $\gamma>1$? It is easy to see that the maximum likelihood (ML) test is not suitable for this problem in general. Consider the following example. Let unif[a,b] denote the uniform distribution over the interval $[a,b]$. Let $p_{1}=\text{unif}[-1,1]$, $p_{2}=\text{unif}[\epsilon,1+\epsilon]$ and $p=\text{unif}[0,1]$. Observe that $H^{2}(p,p_{1})=\frac{\sqrt{2}}{\sqrt{2}-1}$ and $H^{2}(p,p_{2})=\epsilon$. Thus, $p_{2}$ can be arbitrarily closer to $p$ than $p_{1}$. After observing $m$ i.i.d. samples from $p$, the ML test will output $\arg\max\left\{\prod_{i=1}^{m}p_{1}(X_{i}),\prod_{i=1}^{m}p_{2}(X_{i})\right\}$. The test will output $p_{2}$ only if none of the $m$ samples falls in the interval $0,\epsilon]$. The probability of error is equal to $1-(1-\epsilon)^{m}$ which tends to 11 as $m\rightarrow\infty$. Thus, the ML test is not $\gamma$-robust for any $\gamma>1$.
@@ 4 Results: Lower Bound on $\gamma^{*}$
    Let $\mathcal{D}_{1}^{m}$ denote the following product distribution: pick $p$ uniformly at random from the family $\mathcal{D}_{1}$ and draw m i.i.d. samples from pp. Likewise, $\mathcal{D}_{2}^{m}$ denotes the following product distribution: pick $p$ uniformly at random from the family $\mathcal{D}_{2}$ and draw m i.i.d. samples from pp. We show that $TV(\mathcal{D}_{1}^{m},\mathcal{D}_{2}^{m})\leq\frac{1}{3}$.
@@
Let $\mathcal{D}_{1}^{m}$ denote the following product distribution: pick $p$ uniformly at random from the family $\mathcal{D}_{1}$ and draw m i.i.d. samples from pp. Likewise, $\mathcal{D}_{2}^{m}$ denotes the following product distribution: pick $p$ uniformly at random from the family $\mathcal{D}_{2}$ and draw m i.i.d. samples from pp. We now show that $TV(\mathcal{D}_{1}^{m},\mathcal{D}_{1}^{m})$ is small for an appropriate choice of $N_{m}$. We use [7, Lemma 24], which states that two distributions are close (in TV) if there exists an event $E$ such that both distributions put a large mass on $E$ and conditioned on this event the two distributions are close (in TV). We reproduce the proof for completeness.
@@ 5 Some results on testing with respect to symmetric $\chi^{2}$ distance
Thus, $\mathbb{E}[T(X)]\geq 0$ if $\chi^{2}(p,p_{2})\geq\frac{2(\sqrt{2}+1)}{\sqrt{2}-1}\chi^{2}(p,p_{1})$. Likewise, we can show that $\mathbb{E}[T(X)]\leq 0$ if $\chi^{2}(p,p_{1})\geq\frac{2(\sqrt{2}+1)}{\sqrt{2}-1}\chi^{2}(p,p_{2})$. We can get a lower bound of 33 for robust symmetric $\chi^{2}$ testing. The construction of the lower bound is exactly the same (our choice of parameters is $b=1,a_{2}=1,a_{1}\rightarrow\infty$. We only need to show the following. For any distribution $p$ in the perturbed family around $p_{1}$, we have
@@ 6 Alternate Test for the Composite Testing Problem
We now analyze the expected value of $T(X^{n})$ with respect to the unknown distribution pp. By linearity of expectation, it suffices to analyze for $n=1$.

### v2

@@ 1-B Robustness to imperfect modelling
$H_{0}$ $:\text{Pois}(\lambda_{1})$
$H_{1}$ $:\text{Pois}(\lambda_{2})$
@@
$H_{0}$ $:\gamma d(p,p_{1})\leq d(p,p_{2})$
$H_{1}$ $:d(p,p_{1})\geq\gamma d(p,p_{2}).$
If $p$ is not in $H_{0}$ or in $H_{1}$, then the test can make any decision. The error is defined to be the maximum of type-I and type-II errors. We call a test $\gamma$-robust if it uses finitely many samples and keeps the probability of error below some fixed constant $\delta<\frac{1}{2}$ for any choice of $p_{1}$, $p_{2}$, and $p$. Let $\gamma^{*}$ be the smallest $\gamma$ for which a $\gamma$-robust test exists. We call $\gamma^{*}$ the optimal slack factor for the robust testing problem. Previous works have tried to characterize $\gamma^{*}$ for different distance measures $d(.,.)$. The $\ell_{p}$-norm of a function $f:\mathcal{X}\rightarrow\mathbb{R}$ is given by
@@
where $A=\{x^{n}:p_{1}(x^{n})>p_{2}(x^{n})\}$ and $\mu(A)=\frac{1}{n}\sum_{i=1}^{n}1[X_{i}\in A]$, that is, the fraction of observed samples which fall in the set $A$. The lower bound ($\gamma^{*}\geq 3$) was given in the work of Bousquets, Kane, and Moran [7]. The Hellinger distance between $p_{1}$ and $p_{2}$ is given by
@@ 3 Upper Bound on $\gamma^{*}$
What kinds of tests are $\gamma$-robust for some $\gamma>1$? It is easy to see that the maximum likelihood (ML) test is not suitable for this problem in general. Consider the following example. Let unif[a,b] denote the uniform distribution over the interval $[a,b]$. Let $p_{1}=\text{unif}[-1,1]$, $p_{2}=\text{unif}[\epsilon,1+\epsilon]$ and $p=\text{unif}[0,1]$. Observe that $H^{2}(p,p_{1})=\frac{\sqrt{2}}{\sqrt{2}-1}$ and $H^{2}(p,p_{2})=\epsilon$. Thus, $p_{2}$ can be arbitrarily closer to $p$ than $p_{1}$. After observing $m$ i.i.d. samples from $p$, the ML test will output $\arg\max\left\{\prod_{i=1}^{m}p_{1}(X_{i}),\prod_{i=1}^{m}p_{2}(X_{i})\right\}$. The test will output $p_{2}$ only if none of the $m$ samples falls in the interval $0,\epsilon]$. The probability of error is equal to $1-(1-\epsilon)^{m}$ which tends to $1$ as $m\rightarrow\infty$. Thus, the ML test is not $\gamma$-robust for any $\gamma>1$.
@@ 4 Results: Lower Bound on $\gamma^{*}$
    Let $\mathcal{D}_{1}^{m}$ denote the following product distribution: pick $p$ uniformly at random from the family $\mathcal{D}_{1}$ and draw m i.i.d. samples from $p$. Likewise, $\mathcal{D}_{2}^{m}$ denotes the following product distribution: pick $p$ uniformly at random from the family $\mathcal{D}_{2}$ and draw m i.i.d. samples from $p$. We show that $TV(\mathcal{D}_{1}^{m},\mathcal{D}_{2}^{m})\leq\frac{1}{3}$.
@@
Let $\mathcal{D}_{1}^{m}$ denote the following product distribution: pick $p$ uniformly at random from the family $\mathcal{D}_{1}$ and draw m i.i.d. samples from $p$. Likewise, $\mathcal{D}_{2}^{m}$ denotes the following product distribution: pick $p$ uniformly at random from the family $\mathcal{D}_{2}$ and draw m i.i.d. samples from $p$. We now show that $TV(\mathcal{D}_{1}^{m},\mathcal{D}_{1}^{m})$ is small for an appropriate choice of $N_{m}$. We use [7, Lemma 24], which states that two distributions are close (in TV) if there exists an event $E$ such that both distributions put a large mass on $E$ and conditioned on this event the two distributions are close (in TV). We reproduce the proof for completeness.
@@ 5 Some results on testing with respect to symmetric $\chi^{2}$ distance
Thus, $\mathbb{E}[T(X)]\geq 0$ if $\chi^{2}(p,p_{2})\geq\frac{2(\sqrt{2}+1)}{\sqrt{2}-1}\chi^{2}(p,p_{1})$. Likewise, we can show that $\mathbb{E}[T(X)]\leq 0$ if $\chi^{2}(p,p_{1})\geq\frac{2(\sqrt{2}+1)}{\sqrt{2}-1}\chi^{2}(p,p_{2})$. We can get a lower bound of $3$ for robust symmetric $\chi^{2}$ testing. The construction of the lower bound is exactly the same (our choice of parameters is $b=1,a_{2}=1,a_{1}\rightarrow\infty$. We only need to show the following. For any distribution $p$ in the perturbed family around $p_{1}$, we have
@@ 6 Alternate Test for the Composite Testing Problem
We now analyze the expected value of $T(X^{n})$ with respect to the unknown distribution $p$. By linearity of expectation, it suffices to analyze for $n=1$.